# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 6. SPECIAL Planner
A major limitation of existing DPSCA designs is sub-optimal execution planning. Without the ability to pre-estimate intermediate result sizes, traditional methods cannot identify the most efficient execution strategies. To address this, SPECIAL employs an innovative query planner that leverages available synopses for plan size estimation, enabling optimal execution planning.
At a high level, our planner is modeled on the Selinger-style (cost-based) optimizer (Blasgen et al., 1981). It uses a bottom-up, dynamic programming approach to enumerate all equivalent secure execution plans for a given query and estimate their costs to select the optimal one. However, challenges in the design space still remain. First, the introduction of SPECIAL primitives can significantly impact the cost modeling for oblivious operations, rendering existing models (Bater et al., 2018; Liagouris et al., 2023) ineffective. Additionally, the equivalent plan search space is extensive, requiring further strategies to simplify the search space. To address these challenges, we first conduct a comprehensive analysis of the complexities of SPECIAL primitives and develop a new cost model (§ 6.1) for SPECIAL planner. Next, we design protocol-specific heuristics (§ 6.2) tailored for the SPECIAL planner to simplify the plan search space. Note that existing features of the Selinger-style optimizer, such as plan enumerations, are not within our contributions. For detailed understanding of these features, we encourage readers to explore the original IBM System R paper (Blasgen et al., 1981).
@@ 6.1. Cost Model
Similar to the cost model proposed by Shrinkwrap (Bater et al., 2018), we view the cost of a secure execution plan cost as a combination of each operator’s I/O and secure evaluation costs. In addition, we formulate these costs as functions over their processed data sizes. Specifically, given a plan with $\ell$ operators, $\mathsf{op}_{1},...,\mathsf{op}_{\ell}$, and let $\mathbf{I}=\{{I}_{1},...,{I}_{\ell}\}$, $\mathbf{O}=\{{O}_{1},...,{O}_{\ell}\}$, to be the input and output sizes of each operators. The plan cost is formulated as:
@@
Here, $C_{\mathsf{in}}$ represents the data access cost (input I/O), primarily capturing the expenses when moving data from persistent storage to an in-memory secure array. If the accessed data has already loaded into memory, $C_{\mathsf{in}}$ becomes negligible. $C_{\mathsf{out}}$ denotes the output I/O cost, modeling the expenses when writing operator results into output arrays. $C_{\mathsf{eval}}$ accounts for the secure computing expenses for evaluating each operators, typically constituting the dominant cost. Note that, in practice, the exact formulas for $C_{\mathsf{in}},C_{\mathsf{out}}$, and $C_{\mathsf{eval}}$ can vary depending on the specific secure protocol employed (garbled circuits (Yao, 1986), secret sharing (Micali et al., 1987), etc.) as well as the particular hardware configurations in use. Nonetheless, the understanding of the asymptotic costs is adequate for comprehending the principles of our query planning and optimization strategies (Liagouris et al., 2023; Bater et al., 2018). In what follows, we provide detailed analysis on the asymptotic costs for each SPECIAL operator. Similarly, we assume that all input data sizes mentioned henceforth in this section are of size $n$, and all 1D histograms have $m$ bins. Table 1 summarizes the operator costs.
@@
Assuming the maximum size of the indexed buckets is bounded by $O(\frac{n}{\log n})$.
@@
Oblivious sorting and compaction. Oblivious sorting and compaction (OPAC) are crucial for many SPEop operators, so we will begin our analysis with these two primitives. While oblivious sorting algorithms with optimal $O(n\log n)$ complexity exist, they often necessitate either impractically large constants (Ajtai et al., 1983; Goodrich, 2014) or client-side memory (Asharov et al., 2020), both do not fit with SCA scenario. Consequently, we will consider the well-established bitonic sorting based implementation for oblivious sort, which come with $O(n\log^{2}n)$ complexity. Nonetheless, efficient OPAC implementations with $O(n\log n)$ complexity remain achievable (Sasy et al., 2022)
@@
Projection, grouping and aggregation. The PROJECT accesses private relations and discards unnecessary columns independently on each server, without secure computations. Thus, I/O costs dominate this operation, with both input and output costs bounded by $O(n)$. The costs of ORDER-BY, DISTINCT, and GROUP-BY are primarily due to oblivious sorting, resulting in a complexity of $O(n\log^{2}n)$. Additionally, as these operators do not reduce output sizes, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are bounded by $O(n)$. Finally, the cost of aggregations, i.e. COUNT, SUM, and MIN/MAX subjects to a oblivious linear scan, typically outputting a single secret-shared value. Hence, its $C_{\mathsf{eval}}$ is bounded by $O(n)$, with $C_{\mathsf{in}}$ at $O(n)$ and $C_{\mathsf{out}}$ at $O(1)$.
@@
Selections. The primary cost of SELECT stems from an oblivious linear scan, making $C_{\mathsf{eval}}$ within $O(n)$. Since SELECT does not shrink the output size, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are within $O(n)$. (OP)SELECT requires an oblivious compaction (OPAC) before writing outputs, where OPAC usually yields an $O(n\log n)$ complexity (Sasy et al., 2022). Consequently, its $C_{\mathsf{eval}}$ is bounded by $O(n\log n)$ with input I/O cost same as SELECT. However, as (OP)SELECT compacts output size, $C_{\mathsf{out}}$ is reduced to $\mathsf{hist\_bound}=O\left(\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\right)$, where $B_{i}\cap vals\neq\emptyset$ are bins in the synopsis histogram intersecting with selection conditions. If $\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\sim O(1)$, (SP)SELECT becomes preferable, with its running cost dominated by a two-phase linear scan, and thus $C_{\mathsf{eval}}$ is now $O(n)$, and the output cost is $O(1)$. (DC)SELECT is the most efficient selection, though it requires indexable input data. It simply loads indexed data into secure arrays without secure computations or result write-backs. As such, both $C_{\mathsf{evals}}$ and $C_{\mathsf{out}}$ are negligible, with $C_{\mathsf{in}}$ within $\mathsf{idx\_bound}=O\left(\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})-\min_{i}(\mathsf{idx}_{i}.\mathsf{lo})\right)$. Here, $\mathsf{idx}_{i}$ are indexed regions that intersect with selection conditions.
@@
Joins. Both JOIN and (MX)JOIN have $O(n)$ data access costs, but differ in $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$. JOIN, using a Cartesian product approach, doesn’t reduce output size, so both $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$ are within $O(n^{2})$. Compared to JOIN, in the worst-case scenario where the join keys follow a highly biased distribution, i.e. max bucket size reaches $O(n)$, (MX)JOIN’s asymptotic cost is at most $O(n^{2}\log n)$. However, when join keys are distributed more uniformly, the cost can be asymptotically better. For instance, with $m=\log n$ and assuming a max bucket size of $O(\frac{n}{\log n})$, each bucket join costs $O(\frac{n^{2}}{\log n})$, leading to a total cost of $O(n^{2})$, equivalent to JOIN. Recall that bucket joins in (MX)JOIN can be executed concurrently, hence, the processing latency is indeed dominated by the bucket-wise cost, i.e. $O(\frac{n^{2}}{\log n})$. Additionally, the output cost is lowered from $O(n^{2})$ to the sum of per-bucket MF upper bounds (Alg 3:11), which can be substantially less if the join key MFs are relatively low.
@@ 6.2. Heuristics
H-1. Filter push down. Filter pushdown is a widely used optimization technique in conventional databases, where it involves moving the selection operation to the earliest stage in the data processing pipeline, such as directly to the base relation. This reduces the amount of data that needs to be loaded, transferred, and processed, by following operators especially complex joins. However, in SCA designs, data obliviousness typically necessitates padding selection sizes to the worst-case scenario, rendering filter pushdown ineffective (Liagouris et al., 2023; Eskandarian and Zaharia, 2017; Zheng et al., 2017; Poddar et al., 2021). With SPEop selections, the selection outputs can be significantly compacted, making filter pushdown effective again. Therefore, we include filter push down as one of the optimization heuristic for SPEplan.
H-2. Predicates fusion. Let $R$ to be any relation, $A_{1},A_{2},...,A_{k}\subseteq attr(D)$, and $\mathbf{v}=\{v_{1},v_{2},...,v_{k}\}$. We say that for multiple selection over $R$ such that $\sigma_{A_{1}\in v_{1}}(...\sigma_{A_{k}\in v_{k}}(R))$, one can always fuse them into one selection $\sigma_{\mathbf{A}\subset\mathbf{v}}(R)$. This can reduce the number of secure computation invocations from $k$ rounds to just one. Additionally, the selection size can be estimated as $\min_{A_{i}}\left(\mathtt{CardEst}(\sigma_{A_{i}\in v_{i}}(R))\right)$.
@@ 7. Evaluation
In this section, we present evaluation results of our proposed framework. Specifically, we address the following questions:
  • Question-1: Does SPECIAL offer efficiency advantages over current SCA approaches? To what extent can SPECIAL enhance efficiency compared to existing state-of-the-arts (SOTAs)?
  • Question-2: For the DP-based SPECIAL design, is there a trade-off between privacy and efficiency? Can we adjust privacy levels to achieve various efficiency objectives?
  • Question-3: Can SPECIAL scale complex analytical (e.g. multi-way join) queries to large-scale (multi-million rows) datasets?
@@ 7.1. Experimental setups
We briefly outline our experimental setup, covering the baseline systems, prototype SPECIAL implementation, workloads, datasets, and default configurations.
Baseline systems and SPECIAL prototype. We compare SPECIAL with two baseline systems: Shrinkwrap (Bater et al., 2018), the SOTA DPSCA, and SMCQL (Bater et al., 2017), a typical SCA system using exhaustive padding for leakage hiding, also used as a baseline for Shrinkwrap. For consistency, we consider the same circuit-model implementations for both baseline systems and the SPECIAL prototype. While some works (Poddar et al., 2021; Liagouris et al., 2023) similar to SMCQL use exhaustive padding but improve efficiency through protocol-level optimizations, we exclude them from our benchmarks for fair comparison concerns. However, it’s important to note that these optimizations are orthogonal to both the baseline systems and our designs. We re-implement important query features for the two baseline systems, and built SPECIAL using the same MPC package, specifically the sh2pc-0.2.2 protocol within EMPtoolkit-0.2.5. Since managing DP synopses and generating indexes do not involve secure computations, thus we implement the related features using Python (3.7) scripts. All implementations are open-sourced (Lovingmage, 2024).
Datasets and workloads. Although SMCQL and Shrinkwrap use the HealthLNK dataset, it is not public data and thus not accessible to the broader research community. Hence, we developed a new benchmark based on a publicly available dataset, and mimic Shrinkwrap’s workload to create a set of testing queries. Specifically, we use the Czech Financial Dataset (Rel, [n.d.]), an anonymized Czech bank transaction dataset for loan applications. This dataset comprises 8 relational tables with a total of 55 columns and 1.1 million rows. For scaling experiments, we use the raw data and schema to generate synthetic data with up to 10 million rows. Shrinkwrap’s benchmark evaluates four queries: a range query, two 2 binary joins, and a 3-way join. Mirroring this approach, we designed eight query workloads based on the Financial data. All workloads are selection-projection-join (SPJ) queries followed by statistical aggregations such as count, count distinct, and group by. The workloads range from simple linear queries to complex multi-way joins. Due to space limitations, a brief summary of the workloads is provided in Table 2, with the full specification in Appendix B.
@@
Query No. Type Description
Q1 Linear Point query with single predicate.
Q2 Linear Range query with conjunctive predicates.
Q3 Binary Join Non-expanding binary join.
Q4 Binary Join Expanding binary join.
Q5 Multi Join 3-way mixed non-expanding and expanding joins.
Q6 Multi Join 3-way all expanding joins.
Q7 Multi Join 4-way mixed non-expanding and expanding joins.
Q8 Multi Join 5-way mixed non-expanding and expanding joins.
Default configurations. For SPECIAL, each relational table is assigned a fixed privacy budget with $\epsilon=1.5$ and $\delta=0.00005$ to release synopses. This setup ensures that the total privacy loss per table does not exceed $\epsilon=1.5$. For all equal-width histograms generated in SPECIAL, we configure them to have at most 8 bins. For Shrinkwrap, we use a privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ for each query processing, as specified in (Bater et al., 2018). However, Shrinkwrap does not provide bounded privacy loss guarantees, so the privacy guarantee of Shrinkwrap will be no better than SPECIAL. We conduct all experiments on bare-metal Mac machines with M2 Max CPUs and 96GB unified memory. In addition, we consider Q2, Q4, Q8 are default testing queries in each query type, and are used as the representative workload to guide synopses. Moreover, for baseline systems, as they do not have join ordering optimizations, thus we will assume a random join order for them.
Figure 4. In-depth comparisons in execution plans: (i) The exhaustive padding in SMCQL can lead to significant memory blowup; (ii) Both SMCQL and Shrinkwrap suffer from suboptimal join ordering; (iii) Although Shrinkwrap reduces intermediate sizes, it still requires substantial memory to materialize the Cartesian-style standard oblivious join; (iv) SPECIAL effectively identifies optimal join orders and significantly reduces intermediate materialization sizes; (v) The IdxAcc method in SPECIAL not only bypasses secure selections but also directly reduces input I/O costs.
7.2. End-to-end performance comparisons
To address Question-1, we first conduct an end-to-end performance comparison of SPECIAL, Shrinkwrap, and SMCQL across the full benchmark workloads (Q1-Q8). We use two typical performance metrics, query execution time and memory usage in our evaluations. The results are summarized in Figure 5 and Figure 6. Due to significant memory consumption, SMCQL cannot complete the full benchmark and stops at Q4. Therefore, we only include full statistics up to Q3, and for Q4, we provide a projected performance evaluation using 10% of the input data.

### v2

@@ 6. SPECIAL Planner
Current DPSCA designs struggle with costly execution plans because they cannot pre-estimate query intermediate sizes, and thus unable to identify effective execution plans with minimized cost. SPECIAL overcomes this challenge by introducing a novel query planner that uses synopses for size estimation, and thus enabling both private and efficient SCA query planning.
At a high level, our planner is modeled after the Selinger-style optimizer (Blasgen et al., 1981). It uses a bottom-up, dynamic programming approach to enumerate all equivalent plans for a given query, estimates their costs (heavily influenced by intermediate sizes) using available synopses, and selects the plan with minimal cost. The introduction of SPECIAL primitives significantly impacts cost modeling for oblivious operations, rendering existing models (Bater et al., 2018; Liagouris et al., 2023) inadequate. Furthermore, the design-space challenge of query planning persists, and the extensive equivalent plan search space necessitates strategies to simplify the search process. To address these, we first systematically analyze the complexities of SPECIAL primitives and develop a new cost model (§ 6.1). We then design protocol-specific heuristics (§ 6.2) tailored to our planner to narrow the search space.
@@ 6.1. Cost Model
We adopt the standard SCA cost framework (Bater et al., 2018) to develop SPECIAL’s cost model, viewing the cost of a secure execution plan as the sum of each operator’s I/O and secure evaluation costs. Specifically, given a plan with $\ell$ operators, $\mathsf{op}_{1},...,\mathsf{op}_{\ell}$, and let $\mathbf{I}=\{{I}_{1},...,{I}_{\ell}\}$, $\mathbf{O}=\{{O}_{1},...,{O}_{\ell}\}$, to be the input and output sizes of each operators. The plan cost is:
@@
Here, $C_{\mathsf{in}}$ represents the data access cost (input I/O), primarily capturing the expenses when moving data from persistent storage to an in-memory secure array. $C_{\mathsf{out}}$ denotes the output I/O cost, modeling the expenses when writing operator results into output arrays. $C_{\mathsf{eval}}$ accounts for the secure computing cost for evaluating an operator, typically constituting the dominant cost. Note that, in practice, the exact formulas for $C_{\mathsf{in}},C_{\mathsf{out}}$, and $C_{\mathsf{eval}}$ can vary depending on the specific secure protocol employed (garbled circuits (Yao, 1986), secret sharing (Micali et al., 1987), etc.) as well as the particular hardware configurations in use. Nonetheless, the understanding of the asymptotic costs is adequate for comprehending the principles of SCA query planning and optimization strategies (Liagouris et al., 2023; Bater et al., 2018). In what follows, we provide detailed analysis on the asymptotic costs for each SPECIAL operator. Similarly, we assume that all input data sizes mentioned henceforth in this section are of size $n$, and all 1D histograms have $m$ bins. Table 1 summarizes the operator costs.
@@
Assuming the max size of the indexed buckets is bounded by $O(\frac{n}{\log n})$.
@@
Oblivious sorting and compaction. While oblivious sorting algorithms with optimal $O(n\log n)$ complexity exist, they often necessitate either impractically large constants (Ajtai et al., 1983; Goodrich, 2014) or client-side memory (Asharov et al., 2020), both do not fit with SCA scenario. Consequently, we will consider the well-established bitonic sorting based implementation for oblivious sort, which come with $O(n\log^{2}n)$ complexity. Nonetheless, efficient OPAC implementations with $O(n\log n)$ complexity remain achievable (Sasy et al., 2022).
@@
Projection, grouping and aggregation. The PROJECT accesses private relations and discards unnecessary columns independently on each server, which is naturally oblivious. Thus, I/O costs dominate this operation, with both input and output costs bounded by $O(n)$. The costs of ORDER-BY, DISTINCT, and GROUP-BY are dominated by oblivious sorting, resulting in a complexity of $O(n\log^{2}n)$. Additionally, as these operators do not reduce output sizes, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are bounded by $O(n)$. Finally, the cost of aggregations, i.e. COUNT, SUM, and MIN/MAX subjects to a oblivious linear scan, typically outputting a single secret-shared value. Hence, its $C_{\mathsf{eval}}$ is bounded by $O(n)$, with $C_{\mathsf{in}}$ at $O(n)$ and $C_{\mathsf{out}}$ at $O(1)$.
@@
Selections. The primary cost of SELECT stems from an oblivious linear scan, making $C_{\mathsf{eval}}$ within $O(n)$. Since SELECT does not shrink the output size, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are within $O(n)$. (OP)SELECT requires an oblivious compaction (OPAC) before writing outputs, where OPAC usually yields an $O(n\log n)$ complexity (Sasy et al., 2022). Consequently, its $C_{\mathsf{eval}}$ is bounded by $O(n\log n)$ with input I/O cost same as SELECT. However, as (OP)SELECT compacts output size, $C_{\mathsf{out}}$ is reduced to $\mathsf{hist\_bound}=O\left(\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\right)$, where $B_{i}\cap vals\neq\emptyset$ are bins in the synopsis histogram intersecting with selection conditions. If $\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\sim O(1)$, (SP)SELECT becomes preferable, with its running cost dominated by a two-phase linear scan, and thus $C_{\mathsf{eval}}$ is now $O(n)$, and the output cost is $O(1)$. (DC)SELECT is the most efficient selection, though it requires indexable input data. All costs are directly related to the size of the indexed data, so $C_{\mathsf{in}}$, $C_{\mathsf{evals}}$ and $C_{\mathsf{out}}$ are all bounded by $\mathsf{idx\_bound}=O\left(\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})-\min_{i}(\mathsf{idx}_{i}.\mathsf{lo})\right)$. Here, $\mathsf{idx}_{i}$ are indexed regions that intersect with selection conditions.
@@
Joins. Both JOIN and (MX)JOIN have $O(n)$ data access costs, but differ in $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$. JOIN, conducting a Cartesian product for two input tables, has $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$ both bounded by $O(n^{2})$. Compared to JOIN, in the worst-case scenario where the join keys follow a highly biased distribution, i.e. max bucket size reaches $O(n)$, (MX)JOIN’s asymptotic cost is at most $O(n^{2}\log n)$. However, when join keys are distributed more uniformly, the cost can be asymptotically better. For instance, with $m=\log n$ and assuming a max bucket size of $O(\frac{n}{\log n})$, each bucket join costs $O(\frac{n^{2}}{\log n})$, leading to a total cost of $O(n^{2})$, equivalent to JOIN. Recall that bucket joins in (MX)JOIN can be executed concurrently, hence, the processing latency is indeed dominated by the bucket-wise cost, i.e. $O(\frac{n^{2}}{\log n})$. Additionally, the output cost is lowered from $O(n^{2})$ to the sum of per-bucket MF upper bounds (Alg 3:11), which can be substantially less if the join key MFs are low.
@@ 6.2. Heuristics
H-1. Filter push down. is a common query planning optimization, moves selection operations to the earliest possible stage to reduce data processed by subsequent operations. In conventional SCAs, data obliviousness often requires padding selection sizes, making filter pushdown ineffective (Liagouris et al., 2023; Eskandarian and Zaharia, 2017; Zheng et al., 2017; Poddar et al., 2021). However, SPECIAL’s innovative selection methods enable compacting selections to approximate true cardinalities without compromising privacy, restoring the effectiveness of filter pushdown. Hence, we include filter push down as one of the optimization heuristic for SPEplan.
H-2. Predicates fusion. Let $R$ to be any relation, $A_{1},A_{2},...,A_{k}\subseteq attr(R)$, and $\mathbf{v}=\{v_{1},v_{2},...,v_{k}\}$. We say that for multiple selection over $R$ such that $\sigma_{A_{1}\in v_{1}}(...\sigma_{A_{k}\in v_{k}}(R))$, one can always fuse them into one selection $\sigma_{\mathbf{A}\subset\mathbf{v}}(R)$. This can reduce the number of secure computation invocations from $k$ rounds to just one. Additionally, the selection size can be estimated as $\min_{A_{i}}\left(\mathtt{CardEst}(\sigma_{A_{i}\in v_{i}}(R))\right)$.
@@ 7. Evaluation
In this section, we present evaluation results of our proposed framework. Specifically, we address the following questions: Question-1: Does SPECIAL offer efficiency and privacy advantages over existing SCAs? Question-2: For SPECIAL design, is there a trade-off between privacy guarantees and system efficiency? How different synopses scenarios affect system performance? Question-3: Can SPECIAL scale complex analytical (e.g. multi-way join) queries to large-scale (multi-million rows) datasets?
@@ 7.1. Experimental setups
Baseline systems and SPECIAL prototype. We compare SPECIAL with two baseline systems: Shrinkwrap (Bater et al., 2018), the SOTA DPSCA, and SMCQL (Bater et al., 2017) (also used as a baseline for Shrinkwrap). For consistency, we consider the same circuit-model implementations for both baseline systems and the SPECIAL prototype. While some works (Poddar et al., 2021; Liagouris et al., 2023) similar to SMCQL use exhaustive padding but improve efficiency through protocol-level optimizations, we exclude them from our benchmarks for fair comparison concerns. We re-implemented key query features for the two baseline systems and built SPECIAL using the same MPC package ( EMPtoolkit-0.2.5). All implementations are open-sourced (Lovingmage, 2024).
Datasets and workloads. We developed two open benchmarks. The first reproduces the HealthLNK benchmark used by Shrinkwrap and SMCQL which simulates a real-world scenario where medical researchers want to perform secure analytics across multiple cohorts’ sensitive data. We use an open schema (SMCQL, [n.d.]) to generate a scalable synthetic dataset with 4 tables,  [^2] K rows and 24 columns. The benchmark involves four multi-cohort medical study queries, identical to those used in Shrinkwrap and SMCQL. Our second benchmark simulates secure collaborative analytics within the financial sector. Imagine multiple banks needing to analyze their combined, private data to study loan and financial statistics—all without compromising sensitive customer information. We use the anonymized Czech Financial Dataset (Rel, [n.d.]) for this, assuming each entry represents data owned by a different bank or financial organization unable to directly share information. This dataset comprises 8 relational tables with a total of 55 columns and 1.1 million rows. For testing workloads, we design eight query workloads, ranging from simple linear queries to complex multi-way join-aggregation queries. A brief summary of the workloads is provided in Table 2.
Default configurations. For SPECIAL, we employ a per-table privacy budget allocation strategy. Each table is allocated a one-time privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ for synopses generation, and is evenly distributed across all DP synopses. For Shrinkwrap, we adopt their default per-query privacy allocation, assigning a privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ to each query, as outlined in (Bater et al., 2018). It is important to note that this configuration means Shrinkwrap will not offer guarantees on the bounded privacy loss across multiple queries. For the HealthLNK benchmark, we use Dosage and Aspirin as representative workloads, and for the Financial benchmark, we use FQ2, FQ4, and FQ8. Unless further elaborated, these workloads will also serve as default testing queries for our evaluation. For all equal-width histograms generated in SPECIAL, we configure them to have at most 8 bins. Moreover, for baseline systems, as they do not have join ordering optimizations, thus we will assume a random join order for them. We conduct all experiments on bare-metal Mac machines with M2 Max CPUs and 96GB unified memory.
@@
Bench. Query Type Description
HealthLNK Dosage Study Binary Join Expanding binary join.
Comorbidity Binary Join Non-expanding binary join.
Aspirin count Multi Join 3 way mixed join
3 Join Aspirin Multi Join 4 way mixed join
Financial FQ1 Linear Point query.
FQ2 Linear Range query.
FQ3 Binary Join Non-expanding binary join.
FQ4 Binary Join Expanding binary join.
FQ5 Multi Join 3-way mixed joins.
FQ6 Multi Join 3-way all expanding joins.
FQ7 Multi Join 4-way mixed joins.
FQ8 Multi Join 5-way mixed joins.
Figure 3. In-depth comparisons in execution plans: (i) The exhaustive padding in SMCQL can lead to significant memory blowup; (ii) Both SMCQL and Shrinkwrap suffer from unoptimized join ordering; (iii) Although Shrinkwrap reduces intermediate sizes, it still requires substantial memory to materialize join outputs; (iv) SPECIAL can identify efficient join execution orders to reduce intermediate sizes; (v) IdxAcc can significantly reduce input I/O costs.
7.2. End-to-end comparisons
To address Question-1, we first conduct an end-to-end performance comparison of SPECIAL, Shrinkwrap, and SMCQL across all benchmark workloads. The results are summarized in Figure 45. We cannot complete full benchmark for SMCQL due to high memory cost, so we project evaluations for FQ4 (using 10% data) and omit results for other complex workloads.